This is the third time I'm posting this problem, I had two different responses and I am just confused now, please use EXCEL!!!! Thank you!

1 Chapter 1 Problem 14 2 a. Prepare a sources and uses statement for Whistler Corp. for fiscal year 2014 3 b. Prepare a cash flow statement for Whistler Corp. for fiscal year 2014 4 2013 2014 Balance Sheet: Current Assets Cash Accounts Receivable Inventories Total current assets Noncurrent Assets Land Buildings Equipment Patent Accumulated depreciation Total noncurrent assets Total Assets $47,500 $76,700 SO $43,100 $49,000 $36,500 6,500 $156,300 10 12 13 14 15 16 17 18 19 20 S15,800S15,800 S103,600S164,600 S63,200 S65,500 S5,200 S10,800 S12200 238,900 $273,500 S395,200 S5,200 177,000

Current Liabilities Accounts payable to suppliers Income taxes payable Total current liab ilities Noncurrent Liabilities Long tem debt Total liabilities Shareholders Equity Common Stock Retained Earnings Total shareholders equity Total Liabilities and Shareholders Equity 48,000 $25,900 SO S10,700 6,600 23 8,000 25 26 S100,100 $134,000 170,600 $148,100 28 29 S125,400S177,400 SO S47,200 125,400 $224,600 S273,500S395,200 31 32 35 2014 Income statement Sales Cost of goods sold Depreciation expenses Net in terest expenses Other expenses Income before taxes Provision for income taxes Net income Dividends paid Additions to retained eamings 37 $165,300 46,200 1400 -2,000 5,600 110,100 40,900 9,200 22,000 S47,200 39 40 41 45
Answer #1
a) SOURCES AND USES OF FUNDS
SOURCES:
Funds from operations (see statement below) 111500
Long term debt (134000-100100) 33900
Common stock issued (177400-125400) 52000
Total sources 197400
USES:
Dividends 22000
Equipment purchased (65500-63200) 2300
Buildings (164600-103600) 61000
Payment of taxes (40900-10700) 30200
Increase in NWC 81900
Total uses 197400
STATEMENT OF CHANGES IN WORKING CAPITAL
Current Assets: 2013 2014 Increase in NWC Decrease in NWC
Cash 47500 76700 29200
Accounts receivable 0 43100 43100
Inventories 49000 36500 12500
Total current assets 96500 156300
Current liabilities:
Accounts payable to supplliers 48000 25900 22100
Total current liabilities 48000 25900
Net Working Capital 48500 130400
Net increase in working capital 81900 81900
Total 130400 130400 94400 94400
STATEMENT SHOWING FUNDS FROM OPERATIONS
Net income 69200
Add:
Depreciation 1400
Provision for income taxes 40900 42300
Funds from operations 111500
b) Cash flows from operating activities:
Net income 69200
Adjustments to reconcile net income to net cash profived by operations:
Depreciation expense 1400
Increase in accounts receivable = 43100-0 = -43100
Decrease in inventory = 49000-36500 = 12500
Increase in income taxes payable = 10700-0 = 10700
Decrease in accounts payable = 48000-25900 = -22100 -40600
Net cash provided by operating activities 28600
Cash flows from investing activities:
Buildings -61000
Purchase of equipment -2300
Net cash used in investing activities -63300
Cash flows from financing activities:
Issue of common stock 52000
Cash from long term debt 33900
Dividend -22000
Net cash used in financing activities 63900
Net increase (decrease) in cash 29200
Cash balance at beginning of year 47500
Cash balance at end of year 76700